OpenStudy (anonymous):

What is the equation of this line of best fit in slope-intercept form? y=-6x+3/4 y=6x+3/4 y=-3/4x+6 y=3/4x+6

OpenStudy (anonymous):

WILL MEDAL AND FAN @iGreen

OpenStudy (igreen):

What's the y-intercept?

OpenStudy (anonymous):

6?

OpenStudy (igreen):

Yep! y = mx + b Where 'm' is the slope and 'b' is the y-intercept, so we have: y = mx + 6 So our answers can only be C or D.

OpenStudy (anonymous):

is it D?

OpenStudy (igreen):

Now let's find the slope. Take any two points on the line. Let's take (0, 6) and (4, 3) Plug them into the slope-formula: \(\sf m = \dfrac{y_2-y_1}{x_2-x_1}\) \(\sf m = \dfrac{3-6}{4-0}\) Simplify

OpenStudy (anonymous):

CCCC!

OpenStudy (igreen):

Yep! 3 - 6 = -3 4 - 0 = 4 -3/4 So the slope is -3/4 So our final equation is: y = mx + 6 y = -3/4x + 6